OpenStudy (anonymous):

Rewrite the expression so that it does not involve trigonometric functions or inverse trigonometric functions. Csc(sin^-1 x)

OpenStudy (anonymous):

Let \(t=\sin^{-1}x\), so that \(\sin t=x\).  Find the remaining side. Afterwards, you'll have everything you need to find \(\csc t=\csc(\sin^{-1}x)\).

OpenStudy (anonymous):

Yes but I have to rewrite it in terms of not trig functions

OpenStudy (anonymous):

What's the missing side?

OpenStudy (anonymous):

1/x ?

OpenStudy (anonymous):

You have a right triangle you can work with. Use the Pythagorean theorem.

OpenStudy (anonymous):

X-1?